Astronomy Report – November 26th

Astronomy/Science Officer Toluwa Fayemi
As far as our relatively uneventful stay on the Mars Desert Research Station goes, today was a bit more on the eventful side.
The morning started late and we jumped straight into preparing our experiments for the day. Jordan and I spent a good amount of time in the Geology Lab, testing our field 3D scanner on the specimen we retrieved from our EVA the previous day. The results were promising, but we expected to run into some issues in the field. Surely enough, it was difficult to scan an object 360º around while carrying the laptop with us without ‘losing’ the object in the scanner field, particularly in this terrain where the specimen of interest tend to be identical in colour to the ground around it. The sunlight also made it extremely difficult to see the laptop screen, making it difficult to monitor how well we were actually scanning the object.
During the same EVA, we managed to get some exciting photography/videography of the surrounding landscape which may be useful for future exploration/EVAs.
As far as astrophotography, I have been experimenting with me Canon t4i DSLR on various shutter speeds, apertures and ISO settings to continually improve my results. Some of the pictures and the settings used to obtain them are shared below.
This long exposure shot was amongst my first, set at f5.0 and an ISO of 128000, hence the grainy-looking ambient light around the hill. The shutter was open for 30 seconds.
This shot was taken with a briefer shutter, 25seconds and an f-stop of 6.3. This screenshot doesn’t quite do it justice but I’m pleased with the results of this settings modification. I believe the ISO was set to 64000 for this one.
This is the second shot I took with the HAB in view. I set the ISO to 64000 on this one, lower than the 128000 of the previous. The ambient light from the HAB was a tad too much for an ISO so high. This appears to be the better shot. Both shots had a 25 second shutter time and an f-stop of 6.3.

Operations Report – November 26th

Crew 183 Operations Report 26-11-2017

SOL: 4

Name of person filing report: Ashton Stoop

Non-nominal systems:


Notes on non-nominal systems:

Generator (hours run): 14

Solar— 75%

Diesel – 54%

Propane – 40% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 300 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 44

Toilet tank emptied: Yes

ATV’s Used: N/A

Oil Added? No

ATV Fuel Used: 0

# Hours the ATVs were Used today: 0

Notes on ATVs: N/A

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

A lot of the radios have been giving us issues. Sometimes they work and sometimes they don’t. It seems to be a bit random which is a problem when we are trusting them out on 3-hour EVAs. A few of the suits aren’t charging properly. One of the suits failed today during EVA and we had to rush back to the hab before a crew member’s suit was compromised.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well.

Summary of ScienceDome operations:

Sedementation is progressing well.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-There definitely needs to be some further diagnostics performed on suits and radios.

EVA Report – November 26th

EVA #4 Report

The preparation of the 3D scanning device took longer than expected this morning, so we weren’t able to begin depressurization until 1:00 PM. Jordan, Ashton, Toruwa, and I hiked up to the top of Hab Ridge, searching for rocks that had enough detail. We surveyed the area for suitable specimens, but our scans kept running into problems. Since the regolith in the background for the loose boulders have a fairly uniform pattern and color, the scanner had trouble determining what it was supposed to scan. We’ll have to bring our own background with a randomized pattern.

Another issue that we ran into is that the practicality of using the 3D scanner on the field. Since it has to be attached to a laptop via a USB chord, we’d need one person to carry it, and another person to follow close behind while monitoring the status of the 3D scan.

The time that we were out took a surprisingly long time, by the time we descended the hill, it was about 3:00 PM. On the way down, Toruwa lost his footing and fell on his posterior. The fall seemed to deactivate the airflow fan in his suit, meaning we had to get him into the hab fast. By the time we got into the airlock, the inside of his helmet was already fogging up a lot. Since there were four of us, I went into the airlock with Toruwa, while Jordan and Ashton stayed outside, since they needed to further test the 3D scan equipment near the hab.

Due to the circumstances, the Commander ordered an emergency recompression, which took two minutes. Once the recompression was complete, we quickly opened the inner airlock door, getting his helmet off as quickly as we could. He was sweating profusely, and quite exhausted from the ordeal.

Jordan and Ashton’s activities lasted for another 20 minutes, reentering the hab at around 3:25 PM.

We succeeded in achieving what we needed to do on this EVA, but it’s really important to remember how risky it is to go on an EVA, because all it takes is one misstep, and the life support systems could go out.

This is all I have to report.

Thanks Mission Support!

Prepared by Brandon Ferguson

Sol Summary – November 26th

Sol 4 Summary Report

November 26, 2017

Julielynn Wong, Crew 183 Mission Commander

Summary Title: A day of 3D printing, film production, and 3D scanning at MDRS.

Mission Status: Nominal.

Sol Activity Summary:

Today we continued to print 3D4MD digital catalog items out of flexible, rigid and biodegradable plastic feedstock on our humanitarian 3D printing system that can be powered by a portable battery charged by our cycling ergometer system.  

Our EVA team obtained image and video footage and conducted 3D scanning of geological formations with a portable handheld scanner.

Our crew continues to use our cycling ergometer system to follow the exercise countermeasures schedule for astronauts onboard the ISS and to power our humanitarian 3D printing systems. Power generation data was collected non-invasively during training that crew members would have undertaken if no experiment existed.

Our meals were prepared by the Greenhab Officer.

Look Ahead Plan: Preparing for end of sim on Sol 5 at 0700 MST.

Anomalies in work:  None.

Weather:  Calm, sunny and warm.

Crew Physical Status:  Healthy

EVA: Crew Astronomer, Crew Engineer, and Greenhab Officer participated in an EVA led by our Crew Geologist.

Reports to be filed:

GreenHab Report

Operations Report

EVA Report

Astronomer Report

Support Requested: None

GreenHab Report – November 26th

Sol 4 – 26/11/2017

Green Hab Officer Jordan Bowden

Average temperature: 28.5 C in Greenhouse / 8.5 C Outside

Time 11:00 19:00
Temperature inside (°C) 40 C 17 C
Temperature outside (°C) 9C 8 C
% Humidity 18% 21 %

Changes to crops:  Everything is growing well.

Daily water usage for crops: 4 Gallons

Time(s) of watering for crops:  11:00, 17:00

Daily watering and amount of water used:  2 gallons used at 11:00, 2 gallon used at 17:00. 4 gallons total.

Narrative: Greetings CapCom,

Everything was growing well in the Green Hab today – it’s amazing to see the growth since Sol 1. Grow lights are operational. Nothing significant to report.

We had a great EVA today. Progress also continued on our 3D Printing projects – I came close to completing my design for a Neonatal Aspirator device.

All the best,

Jordan Bowden
Green Hab Officer

Copyright © The Mars Society. All rights reserved. | Main Site